OpenStudy (anonymous):

For expressions with some holes, why are they equal to expressions with less holes. For instance, (example is in the comments) Both expressions share a hole of x = 0, but the second one has a hole of x = -1 while the other one does not. I appreciate all insight.

OpenStudy (anonymous):

\[\frac{ 1 }{ x } = \frac{ x+1 }{ x(x+1) }\]

OpenStudy (anonymous):

The two expressions aren't equal to each other for all values of \(x\). When \(x=-1\), you have \[-1=\frac{0}{0}\] which is indeterminate, and when \(x=0\) you have \[\frac{1}{0}=\frac{1}{0}\] which is undefined. But for any other value of \(x\), say \(x=4\), the expressions are equivalent. \[\frac{1}{4}=\frac{5}{20}\]

OpenStudy (anonymous):

The reason the expressions are equivalent for all \((-\infty<x<-1)\cup(-1<x<0)\cup(0<x<\infty)\) is because the discontinuity at \(x=-1\) can be removed; a factor of \(x+1\) is present in both the numerator and denominator, so you can cancel (but only when \(x=-1\); otherwise you cannot).

OpenStudy (anonymous):

Okay, so effect, we are saying they are equal because the hole can be removed. But if we looked at the hole for any other reason, then it would be not equal.